# Scalar-induced gravitational waves from coherent initial states

### v1

@@ V Chirality in secondary GW
We shall begin by inspecting the chirality of the one-point function of the induced tensor perturbations $\braket{h^{\lambda}_{\bm{k}}}$. Let us retain a generic structure of the $\alpha({\bm{k}})$ as given in Eq. (47). We shall not assume any specific form of $F(k)$ and allow for a general angular dependence in the anisotropy inducing factor $\beta(\hat{\bm{k}})$, expanding it in terms of spherical harmonics as
$\beta(\bm{k})$ $=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\beta_{\ell m}Y_{\ell m}(\hat{\bm{k}}).$ (53bh)
@@
×$\times(-1)^{m}\sqrt{(2\ell+1)(2\ell^{\prime}+1)}\sqrt{\frac{(\ell-m)!(\ell^{\prime}-m^{\prime})!}{(\ell+m)!(\ell^{\prime}+m^{\prime})!}}$
×$\times\int{\rm d}q\,q\int{\rm d}(\cos\theta)\frac{\sqrt{\mathcal{P}_{{}_{\mathrm{S}}}(p)\mathcal{P}_{{}_{\mathrm{S}}}(q)}}{p^{3}}F(p)F(q)$
×$\times\sin^{2}\theta P_{\ell m}(\cos\theta)P_{\ell^{\prime}m^{\prime}}\left(\frac{k-q\cos\theta}{p}\right)I(p,q,\eta),$ (53bi)
@@
$m^{\prime}=\begin{cases}-(m+2)&\text{for }L\,,\\ -(m-2)&\text{for }R\,,\end{cases}$ (53bj)
and $\bm{p}\equiv\bm{k}-\bm{q}$. It is evident from the above selection rule that for a constant $\beta(\hat{\bm{k}})$, i.e., where there are only monopole contributions ($\ell,\ell^{\prime}=0$) in $\beta_{\ell m}$, the tensor one-point function identically vanishes for both left- and right-handed modes. Therefore, the primordial mean of the scalar perturbation must violate statistical isotropy to induce a non-zero mean $\braket{h^{\lambda}_{\bm{k}}}$ for the secondary tensor perturbations.
For a general $\beta(\hat{{\bm{k}}})$, $\braket{h^{\lambda}_{\bm{k}}}$ of different polarization modes $\lambda$ are different essentially due to the selection rules differing as $m^{\prime}=m\pm 2$. So, we have a parity-odd one point function as $\braket{h_{\bm{k}}^{L}(\eta)}\neq\braket{h_{\bm{k}}^{R}(\eta)}.$
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}})$ $=$ $\frac{k^{3}}{2\pi^{2}}\,\big{[}f^{\lambda_{1}}(k)\delta^{\lambda_{1}\lambda_{2}}+k^{3}\Delta^{3}g^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}})\big{]}$ (53bk)
@@
×$\times|Q^{L,R}(\bm{k},\bm{q})|^{2}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|\bm{k}-\bm{q}|)}{|\bm{k}-\bm{q}|^{3}}$
[…]
@@
$\mathscr{P}_{{}_{\rm h}}^{LL}({\bm{k}},-{\bm{k}})-\mathscr{P}_{{}_{\rm h}}^{RR}({\bm{k}},-{\bm{k}})=$ $\Delta^{3}\frac{k^{6}}{2\pi^{2}}\big{[}|\braket{h^{L}_{\bm{k}}}|^{2}-|\braket{h^{R}_{\bm{k}}}|^{2}\big{]}\,,$ (53bm)
and it is non-zero due to Eq. (53bi). We may quantify chirality in terms of spectral density, focussing on terms with ${\bm{k}}_{2}=-{\bm{k}}_{1}$ as
@@
$\Pi_{k}$ $\simeq$ $\frac{k^{3}\Delta^{3}[|\braket{h^{L}_{\bm{k}}}|^{2}-|\braket{h^{R}_{\bm{k}}}|^{2}]}{f^{L}(k)+f^{R}(k)}\,.$ (53bp)
@@
$\mathscr{P}_{{}_{\rm h}}^{LR}({\bm{k}},-{\bm{k}})$ $=$ $\Delta^{3}\frac{k^{6}}{2\pi^{2}}\big{[}g^{LR}({\bm{k}},-{\bm{k}})+\braket{h^{L}_{\bm{k}}}\braket{h^{R}_{\bm{k}}}^{\ast}\big{]}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ (53bq)
$=$ $[\mathscr{P}_{{}_{\rm h}}^{RL}({\bm{k}},-{\bm{k}})]^{\ast}\,,$ (53br)
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ $=$ $\left(\frac{k_{1}^{3}+k_{2}^{3}}{2}\right)\Delta^{3}\,\frac{(k_{1}k_{2})^{3/2}}{2\pi^{2}}\big{[}g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ (53bs)
@@
Evidently, the left- and right-circular components in this configuration are different $\mathscr{P}_{{}_{\mathrm{h}}}^{LL}(\bm{k}_{1},\bm{k}_{2})\neq\mathscr{P}_{{}_{\mathrm{h}}}^{RR}(\bm{k}_{1},\bm{k}_{2})$ and once again, we have non-vanishing correlation between different polarization modes $\mathscr{P}_{{}_{\mathrm{h}}}^{LR}(\bm{k}_{1},\bm{k}_{2})\neq 0$.
@@ VI Conclusion
In light of the signature of stochastic GW background in PTA [18, 20, 21], and proposed sensitivities of upcoming missions like SKA [10, 69] and interferometers such as LISA [70, 15, 71], it becomes compelling and feasible to examine the scenario of violation of statistical homogeneity, isotropy and parity using the cosmological GW background. The amplitude and shape of $\Omega_{{}_{\rm GW}}$ around its maximum, if compared against a model of $\alpha(k)$, can inform us about the level of inhomogeneity of the primordial mean, or at least provide an upper bound on it. The chirality of $\Omega_{{}_{\rm GW}}$, if measured in future, shall provide insights into the anisotropic structure of primordial mean induced in the tensor perturbations. The existence of correlation between different modes ${\bm{k}}_{2}\neq{\bm{k}}_{1}$ and/or correlation between different polarizations of $L$ and $R$ (or $+$ and $\times$), shall break degeneracies between this scenario and other possibilities that may lead to enhanced tensor power over small scales. Further, the structure of anisotropies in the $\Omega_{{}_{\rm GW}}$ can be readily computed in terms of $\beta_{\ell m}$ that captures the anisotropy present in $\alpha({\bm{k}})$ in a fairly generic manner [cf. Eqs. (47) and (53bh)]. Such an exercise, especially for modes with $\ell\neq\ell^{\prime}$, shall provide richer information regarding the precise nature of $\alpha({\bm{k}})$. We relegate such an analysis to a future work.
@@ Appendix A Identifying spectral density in case of inhomogeneities
We shall outline an alternative method to compute spectral density for a function, in case of statistical inhomogeneity leading to unavoidable spatial dependence. We shall illustrate it with the energy density of the scalar-induced GW $\rho_{\rm GW}$.
@@
$+\braket{h^{\lambda_{1}}_{\bm{k}_{1}}}\braket{h^{\lambda_{2}}_{\bm{k}_{2}}}\,.$ (53bu)
@@
$f(\eta)$ $=$ $\iint\frac{{\rm d}^{3}{\bm{k}}_{1}{\rm d}^{3}{\bm{k}}_{2}}{(2\pi)^{3}}f(k_{1})\,\delta({\bm{k}}_{1}+{\bm{k}}_{2})\,e^{i\,{\bm{x}}\cdot({\bm{k}}_{1}+{\bm{k}}_{2})}\,.$ (53bv)
Under suitable change of integration variables such that it reduces to
$f(\eta)$ $=$ $\int\frac{{\rm d}^{3}{\bm{k}}}{(2\pi)^{3}}f(k)=\int{\rm d}\ln k\frac{k^{3}}{2\pi^{2}}f(k)\,.$ (53bw)
@@
$g(\eta,{\bm{x}})$ $=$ $\iint\frac{{\rm d}^{3}{\bm{k}}_{1}{\rm d}^{3}{\bm{k}}_{2}}{(2\pi)^{3}}g({\bm{k}}_{1},{\bm{k}}_{2})\,e^{i\,{\bm{x}}\cdot({\bm{k}}_{1}+{\bm{k}}_{2})}\,.$ (53by)
@@
$\frac{{\rm d}g(\eta,{\bm{x}})}{{\rm d}\ln k}=\frac{k^{3}}{(2\pi)^{3}}\int{\rm d}^{2}\Omega_{\hat{\bm{k}}}\,e^{i{\bm{x}}\cdot{\bm{k}}}\int{{\rm d}^{3}{\bm{q}}}\,g({\bm{q}},{\bm{k}}-{\bm{q}})\,.$ (53cb)
@@
$h^{2}(\eta,{\bm{x}})$ $=$ $\bigg{[}\int\frac{{\rm d}^{3}{\bm{k}}}{(2\pi)^{3/2}}h({\bm{k}})\,e^{i\,{\bm{x}}\cdot{\bm{k}}}\bigg{]}^{2}\,,$ (53cc)
$=$ $\iint\frac{{\rm d}^{3}{\bm{k}}\,{\rm d}^{3}{\bm{q}}}{(2\pi)^{3}}h({\bm{q}})h({\bm{k}}-{\bm{q}})\,e^{i{\bm{x}}\cdot{\bm{k}}}\,,$ (53cd)
@@
$\frac{{\rm d}h^{2}(\eta,{\bm{x}})}{{\rm d}\ln k}=\frac{k^{3}}{(2\pi)^{3}}\int{\rm d}^{2}\Omega_{\hat{\bm{k}}}\,e^{i{\bm{x}}\cdot{\bm{k}}}\int{{\rm d}^{3}{\bm{q}}}\,h({\bm{q}})h({\bm{k}}-{\bm{q}})\,.$ (53cg)
Adding Eqs. (53bx) (53cb) and (53cg), we get the spectral density associated with $\rho_{\rm GW}(\eta,{\bm{x}})$ to be

### v2

@@ IV.1 Parameterization of the primordial mean
IV.2 Parameterization of the primordial mean: Dirac $\delta$ function
As another example, we obtain the spectral density of the SIGW using a Dirac $\delta$ function for the scalar primordial mean. This exercise will reveal the analytical form of the enhancement in the SIGW signal due to the statistical inhomogeneities.
Let us consider the inhomogeneous part of the polarization-summed $\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}(\bm{k},-\bm{k}{,\eta})$ [cf. second term of Eq.(39)]
$\sum_{\lambda}{\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}}_{\rm(ih)}(\bm{k},-\bm{k}{,\eta})=$ $\frac{2^{4}}{\pi}k^{3}\int{\rm d}^{3}\bm{q}\sum_{\lambda}Q_{\lambda}^{2}(\bm{k},\bm{q})$
$\times$ $I^{2}(|\bm{k}-\bm{q}|,q,\eta)\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|\bm{k}-\bm{q}|)}{|\bm{k}-\bm{q}|^{3}}$
$\times$ $\Delta^{3}k^{3}\mathinner{\!\left\lvert\alpha(\bm{q})\right\rvert}^{2},$ (60)
where we have performed the transformation $\bm{q}\to\bm{k}-\bm{q}$, and used the symmetries of $Q_{\lambda}(\bm{k},\bm{q})$ and $I(p,q,\eta)$ for further simplifications. Let us consider again the lognormal parametrization of $\alpha({\bm{k}})$ [cf. Eq. (48)] and inspect the functional form of $F^{2}(k)$ appearing in the integrand of $\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}({\bm{k}},-{\bm{k}},\eta)$. We take the limit of $\Delta_{f}\to 0$ which reduces $F^{2}(k)$ to be
$F^{2}(k)$ $=\delta\left(\ln\frac{k}{k_{0}}\right)$ (61)
We proceed as before and use the notations introduced in the previous subsection [cf. Eq. (50)] to write
$\sum_{\lambda}{\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}}_{\rm(ih)}(\bm{k},-\bm{k}{,\eta})$ $=2^{4}\alpha_{0}^{2}\Delta^{3}\frac{k_{0}}{k}\int_{-1}^{1}{\rm d}\mu(1-{\mu}^{2})^{2}\frac{1}{u_{0}^{3}}$
$\times\mathcal{P}_{s}(u_{0}k)\mathcal{P}_{s}(k_{0})\tilde{I}^{2}\left(u_{0},v_{0},x\right)$ (62)
where $v_{0}\equiv k_{0}/k,\ u_{0}\equiv u(v_{0},\mu)=\sqrt{1+v_{0}^{2}-2\mu v_{0}}.$ We can perform the integration analytically in the asymptotic large- and small-scale limits, i.e., when $v_{0}=k_{0}/k\gg 1$ and $k_{0}/k\ll 1$, which will reveal the analytical profiles of the rise and fall of the enhancement in the spectral density. We consider $\mathcal{P}_{{}_{\mathrm{S}}}(k)=A_{\rm s}$ as before.
First consider the regime where $v_{0}\gg 1\Rightarrow k\ll k_{0}$, which implies $u_{0}\approx v_{0}$. The function $\tilde{I}(u_{0},v_{0},x)$ can be approximated in this region as
$\tilde{I}^{2}(v_{0},v_{0},x)\approx\frac{18}{k^{2}\eta^{2}v_{0}^{4}}\ln^{2}v_{0}\,$, (63)
[cf. Eq. (52)]. This leads to the asymptotic expression
$\sum_{\lambda}{\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}}_{\rm(ih)}(\bm{k},-\bm{k}{,\eta})\approx$ $\frac{1536}{5k^{2}\eta^{2}}\alpha_{0}^{2}\Delta^{3}A_{\rm s}^{2}\left(\frac{k}{k_{0}}\right)^{6}\ln^{2}\left(\frac{k}{k_{0}}\right),$
for $k\ll k_{0}$. In the regime where $v_{0}\ll 1\Rightarrow k\gg k_{0}$ (and so $u_{0}\simeq 1$), the function $\tilde{I}(u_{0},v_{0},k)$ can be approximated as $\tilde{I}^{2}(u_{0},v_{0},k)\approx 2/{k^{2}\eta^{2}}$, leading to the asymptotic form of
$\sum_{\lambda}{\mathscr{P}_{{}_{\rm h}}^{\lambda\lambda}}_{\rm(ih)}(\bm{k},-\bm{k}{,\eta})\approx$ $\frac{2^{5}}{k^{2}\eta^{2}}\alpha_{0}^{2}\Delta^{3}A_{\rm s}^{2}\frac{k_{0}}{k}\,$, (65)
for $k\gg k_{0}$. Using these expressions, we obtain the asymptotic behaviors of the diagonal part of the spectral density $\Omega_{{}_{\rm GW}}^{0}(\bm{k},-\bm{k})$ [cf. Eq. (46)]. Figure 3 shows $\Omega_{{}_{\rm GW}}^{0}(\bm{k},-\bm{k})$ obtained from the exact numerical solution Eq. (62) (black plot), the analytical solution for small $k$ Eq. (LABEL:eq:delta_Ph_small_k) (green plot), and the analytical solution for large $k$ Eq. (65) (blue plot). As the analytical results indicate, the enhancement in the spectral density has a sharp rise $\Omega_{{}_{\rm GW}}^{0}\propto(k/k_{0})^{6}\ln^{2}(k/k_{0})$, followed by a more gradual decline $\Omega_{{}_{\rm GW}}^{0}\propto(k/k_{0})^{-1}$ after peaking around $k/k_{0}=1$.
Figure 3: Enhancement in the spectral density of secondary gravitational waves induced by a scalar primordial mean with a Dirac-$\delta$ profile (cf. Eq. (61)). The black curve shows inhomogeneous contribution to $\Omega_{{}_{\rm GW}}^{0}(\bm{k},-\bm{k})$ obtained from the exact numerical solution Eq. (62), while the green and blue curves show the analytical solutions for small and large $k$ limits (Eq. (LABEL:eq:delta_Ph_small_k) and Eq. (65)), respectively. The enhancement in the spectral density rises $\propto(k/k_{0})^{6}\ln^{2}(k/k_{0})$, followed by a more gradual fall $\propto(k/k_{0})^{-1}$ after peaking near $k/k_{0}=1$. We have set the model parameters to be $\mathcal{P}_{{}_{\mathrm{S}}}(k)=A_{\rm s}=10^{-9}$, $\alpha_{0}=0.1/\sqrt{A_{\rm s}}$, and $\Delta=0.1$.
@@ V Chirality in secondary GW
[…]
@@
$\times(-1)^{m}\sqrt{(2\ell+1)(2\ell^{\prime}+1)}\sqrt{\frac{(\ell-m)!(\ell^{\prime}-m^{\prime})!}{(\ell+m)!(\ell^{\prime}+m^{\prime})!}}$
$\times\int{\rm d}q\,q\int{\rm d}(\cos\theta)\frac{\sqrt{\mathcal{P}_{{}_{\mathrm{S}}}(p)\mathcal{P}_{{}_{\mathrm{S}}}(q)}}{p^{3}}F(p)F(q)$
$\times\sin^{2}\theta P_{\ell m}(\cos\theta)P_{\ell^{\prime}m^{\prime}}\left(\frac{k-q\cos\theta}{p}\right)I(p,q,\eta),$ (67)
@@
[…]
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}})$ $=$ $\frac{k^{3}}{2\pi^{2}}\,\big[f^{\lambda_{1}}(k)\delta^{\lambda_{1}\lambda_{2}}+k^{3}\Delta^{3}g^{\lambda_{1}\lambda_{2}}({\bm{k}},-{\bm{k}})\big]$ (69)
@@
$\times|Q^{L,R}(\bm{k},\bm{q})|^{2}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|\bm{k}-\bm{q}|)}{|\bm{k}-\bm{q}|^{3}}$
$\times\mathinner{\!\left\lvert\alpha(\bm{k}-\bm{q})\right\rvert}^{2}\,$, (70)
@@
$\mathscr{P}_{{}_{\rm h}}^{LL}({\bm{k}},-{\bm{k}})-\mathscr{P}_{{}_{\rm h}}^{RR}({\bm{k}},-{\bm{k}})=$ $\Delta^{3}\frac{k^{6}}{2\pi^{2}}\big[|\braket{h^{L}_{\bm{k}}}|^{2}-|\braket{h^{R}_{\bm{k}}}|^{2}\big]\,$, (71)
and it is non-zero due to Eq. (67). We may quantify chirality in terms of spectral density, focussing on terms with ${\bm{k}}_{2}=-{\bm{k}}_{1}$ as
@@
$\Pi_{k}$ $\simeq$ $\frac{k^{3}\Delta^{3}[|\braket{h^{L}_{\bm{k}}}|^{2}-|\braket{h^{R}_{\bm{k}}}|^{2}]}{f^{L}(k)+f^{R}(k)}\,.$ (74)
@@
$\mathscr{P}_{{}_{\rm h}}^{LR}({\bm{k}},-{\bm{k}})$ $=$ $\Delta^{3}\frac{k^{6}}{2\pi^{2}}\big[g^{LR}({\bm{k}},-{\bm{k}})+\braket{h^{L}_{\bm{k}}}\braket{h^{R}_{\bm{k}}}^{\ast}\big]\,$$penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(75)
$=$ $[\mathscr{P}_{{}_{\rm h}}^{RL}({\bm{k}},-{\bm{k}})]^{\ast}\,$, (76)
@@
$\mathscr{P}_{{}_{\rm h}}^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ $=$ $\left(\frac{k_{1}^{3}+k_{2}^{3}}{2}\right)\Delta^{3}\,\frac{(k_{1}k_{2})^{3/2}}{2\pi^{2}}\big[g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$ (77)
@@
Evidently, the left- and right-circular components in this configuration are different $\mathscr{P}_{{}_{\mathrm{h}}}^{LL}(\bm{k}_{1},\bm{k}_{2})\neq\mathscr{P}_{{}_{\mathrm{h}}}^{RR}(\bm{k}_{1},\bm{k}_{2})$ and once again, we have non-vanishing correlation between different polarization modes $\mathscr{P}_{{}_{\mathrm{h}}}^{LR}(\bm{k}_{1},\bm{k}_{2})\neq 0$. We should point out that the non-vanishing values of correlations between different polarizations (such as $L$ and $R$ or equivalently ++ and $\times$) is a signature of parity-violation occurring along with the violation of homogeneity and isotropy. Hence, in case of degeneracy between our scenario of interest and other mechanisms of parity-violation leading to similar $\Pi_{k}$, such a correlation between polarizations can resolve the degeneracy.
@@ VI Conclusion
In light of the signature of stochastic GW background in PTA [20, 22, 23], and proposed sensitivities of upcoming missions like SKA [10, 74] and interferometers such as LISA [75, 17, 76], it becomes compelling and feasible to examine the scenario of violation of statistical homogeneity, isotropy and parity using the cosmological GW background. The amplitude and shape of $\Omega_{{}_{\rm GW}}$ around its maximum, if compared against a model of $\alpha(k)$, can inform us about the level of inhomogeneity of the primordial mean, or at least provide an upper bound on it. The chirality of $\Omega_{{}_{\rm GW}}$, if measured in future, shall provide insights into the anisotropic structure of primordial mean induced in the tensor perturbations. The existence of correlation between different modes ${\bm{k}}_{2}\neq{\bm{k}}_{1}$ and/or correlation between different polarizations of $L$ and $R$ (or ++ and $\times$), shall break degeneracies between this scenario and other possibilities that may lead to enhanced tensor power over small scales. Further, the structure of anisotropies in the $\Omega_{{}_{\rm GW}}$ can be readily computed in terms of $\beta_{\ell m}$ that captures the anisotropy present in $\alpha({\bm{k}})$ in a fairly generic manner [cf. Eqs. (47) and (66)]. Such an exercise, especially for modes with $\ell\neq\ell^{\prime}$, shall provide richer information regarding the precise nature of $\alpha({\bm{k}})$. We relegate such an analysis to a future work.
@@ Appendix A Identifying spectral density in case of inhomogeneities
We shall outline an alternative method to compute spectral density for a function, in case of statistical inhomogeneity leading to unavoidable spatial dependence. We shall illustrate it with the energy density of the SIGW $\rho_{\rm GW}$.
@@
$+\braket{h^{\lambda_{1}}_{\bm{k}_{1}}}\braket{h^{\lambda_{2}}_{\bm{k}_{2}}}\,.$ (79)
@@
$f(\eta)$ $=$ $\iint\frac{{\rm d}^{3}{\bm{k}}_{1}{\rm d}^{3}{\bm{k}}_{2}}{(2\pi)^{3}}f(k_{1})\,\delta({\bm{k}}_{1}+{\bm{k}}_{2})\,e^{i\,{\bm{x}}\cdot({\bm{k}}_{1}+{\bm{k}}_{2})}\,.$ (80)
Under suitable change of integration variables, it reduces to
$f(\eta)$ $=$ $\int\frac{{\rm d}^{3}{\bm{k}}}{(2\pi)^{3}}f(k)=\int{\rm d}\ln k\frac{k^{3}}{2\pi^{2}}f(k)\,.$ (81)
@@
$g(\eta,{\bm{x}})$ $=$ $\iint\frac{{\rm d}^{3}{\bm{k}}_{1}{\rm d}^{3}{\bm{k}}_{2}}{(2\pi)^{3}}g({\bm{k}}_{1},{\bm{k}}_{2})\,e^{i\,{\bm{x}}\cdot({\bm{k}}_{1}+{\bm{k}}_{2})}\,.$ (83)
@@
$\frac{{\rm d}g(\eta,{\bm{x}})}{{\rm d}\ln k}=\frac{k^{3}}{(2\pi)^{3}}\int{\rm d}^{2}\Omega_{\hat{\bm{k}}}\,e^{i{\bm{x}}\cdot{\bm{k}}}\int{{\rm d}^{3}{\bm{q}}}\,g({\bm{q}},{\bm{k}}-{\bm{q}})\,.$ (86)
@@
$h^{2}(\eta,{\bm{x}})$ $=$ $\bigg[\int\frac{{\rm d}^{3}{\bm{k}}}{(2\pi)^{3/2}}h({\bm{k}})\,e^{i\,{\bm{x}}\cdot{\bm{k}}}\bigg]^{2}\,$, (87)
$=$ $\iint\frac{{\rm d}^{3}{\bm{k}}\,{\rm d}^{3}{\bm{q}}}{(2\pi)^{3}}h({\bm{q}})h({\bm{k}}-{\bm{q}})\,e^{i{\bm{x}}\cdot{\bm{k}}}\,$, (88)
@@
$\frac{{\rm d}h^{2}(\eta,{\bm{x}})}{{\rm d}\ln k}=\frac{k^{3}}{(2\pi)^{3}}\int{\rm d}^{2}\Omega_{\hat{\bm{k}}}\,e^{i{\bm{x}}\cdot{\bm{k}}}\int{{\rm d}^{3}{\bm{q}}}\,h({\bm{q}})h({\bm{k}}-{\bm{q}})\,.$ (91)
Adding Eqs. (82) (86) and (91), we get the spectral density associated with $\rho_{\rm GW}(\eta,{\bm{x}})$ to be